OpenStudy (kainui):

What's the most parallel as possible line we can make to the side of the parabola y=x^2 from x=0 to x=infinity?

OpenStudy (zzr0ck3r):

what?

OpenStudy (kainui):



OpenStudy (zzr0ck3r):

you need to pick a point for it to be parallel to, then its the tangent line.

OpenStudy (kainui):

I want it to be most close to being average to the slope at every point.

OpenStudy (zzr0ck3r):

there is parallel and not parallel, there is nothing inbetween

OpenStudy (zzr0ck3r):

my guess is that would be at 0

OpenStudy (zzr0ck3r):

no idea ... lol

OpenStudy (kainui):

Well the slope is different at every point on y=x^2. I want to make the difference between the slope on a line to be as close as possible to the slope at every point. Since the slope only gets more and more positive the further to the right you go, I think a fairly safe first guess is a slope of 1 if that makes sense.

OpenStudy (zzr0ck3r):

but x^2 is always changing, so I dont see how you would compute this "average".

OpenStudy (zzr0ck3r):

ill stay posted to see what others do. interesting:)

OpenStudy (anonymous):

It almost sounds like needing to take the limit of the derivative of y=x^2 ... But that would just be infinity... So maybe make a table of several large numbered x values with their instantaneous slopes and finding a slope that they appear to be approaching? But avoiding all the math, my best guess would be a slope of m=infinity

OpenStudy (anonymous):

If x is infinity and the function is y=x^2 it will never converge, it has no limit, it is divergent. Or u might even say that the slope is undefined because the slope from one point to the next essentially would become vertical.

OpenStudy (anonymous):

draw a slope graph and find the distance vertically